OpenStudy (anonymous):

How would i solve x/6 - y/5 = 1?

OpenStudy (saifoo.khan):

Which variable do you have to sole for?

OpenStudy (saifoo.khan):

solve

OpenStudy (anonymous):

I have to graph it, so I'm guessing y

OpenStudy (saifoo.khan):

For graphing use this: When y-intercepts, make x =0 When x-intercepts, make y =0

OpenStudy (anonymous):

so, when using the y-inercepts, i'll make x 0?

OpenStudy (anonymous):

x = 0?

OpenStudy (saifoo.khan):

x/6 - y/5 = 1 (0)/6 - y/5 = 1 Solve for y. This will give you the y-intercept.

OpenStudy (anonymous):

oh, okay. thanks